OpenStudy (anonymous):

a barrel contains 153 gallons of water and is being drained at a constant rate of 10 gallons per minute write an equation that models the number of gallons after t minutes i have to show my work on this one

OpenStudy (anonymous):

similar to the previous one but here, they actually give you the rate: -10

OpenStudy (anonymous):

so if it starts at 153 gallons and decreases by 10 gallons every minute... what's the equation going to look like?

OpenStudy (anonymous):

g=t-10

OpenStudy (anonymous):

153=t-10

OpenStudy (anonymous):

t= 0 minutes gallons = 153 t=1 minutes gallons = 143 t=2 minutes gallons = 133 

OpenStudy (anonymous):

equations of lines are like: y = rate*x +y intercept y intercept is the amount of whatever we're dealing with when x=0 for acres it was 5 for gallons it's 153

OpenStudy (anonymous):

so y = rate*x +153 ... got that part?

OpenStudy (anonymous):

now put in the rate they give you

OpenStudy (anonymous):

y=...?

OpenStudy (anonymous):

what's the rate

OpenStudy (anonymous):

g=153-10t

OpenStudy (anonymous):

can you please continue this one